1. an article in the transactions of the institution of chemical engineers reported data from an experiment investigating the effect of several process variables on the vapor phase oxidation of naphthalene. A sample of the percentage mole conversion of naphthalene to maleic anhydryde follows 4.2,4.7,4.7,5.0,3.8,3.6,3.0,5.1,3.1,3.8,4.8,4.0,5.2,4.3,2.8,2.0,2.8,3.3,4.8 and 5.0

a calculate the sample mean

b find the medium

c find the mode

d calculate the standard deviation
Answer #1

a) x1= 4.2,x2=4.7,x3=4.7,x4=5.0,x5=3.8,   22.4

x6=3.6,x7=3.0,x8=5.1,x9=3.1,x10=3.8, 18.6

x11=4.8,x12=4.0,x13=5.2,x14=4.3,x15=2.8, 21.1

x16=2.0,x17=2.8,x18=3.3,x19=4.8 ,x20=5.0   17.9

sample mean = (x1 + x2 + .......+ xn ) / n

n= total number of samples = 20

sample mean( \mu ) =(4.2+4.7+4.7+5.0+3.8+3.6+3.0+5.1+3.1+3.8+4.8+4.0+5.2+4.3+2.8+2.0+2.8+3.3+4.8 + 5.0 ) / 20

                     \mu = 4

b) medium :

first arrange the values from minimum to maximum value

[2, 2.8, 2.8, 3, 3.1 , 3.3, 3.6, 3.8 ,3.8, 4 , 4.2, 4.3, 4.7, 4.7, 4.8, 4.8, 5 , 5 , 5.1, 5.2 ]

If n is an even number of data values the median is the average of the two values present in the middle when numbers are in increasing order.

total 20 values are present

here n=20 even number

find the average of 10th and 11 th term in the order.

10th term = 4   11th term=4.2

medium = (4+ 4.2 ) / 2 = 4.1

medium   = 4.1

c) Mode :

Mode is the number which is present in data and that comes maximum number of times.

[2, 2.8, 2.8, 3, 3.1 , 3.3, 3.6, 3.8 ,3.8, 4 , 4.2, 4.3, 4.7, 4.7, 4.8, 4.8, 5 , 5 , 5.1, 5.2 ]

In our case some numbers came two times.So all these numbers considered as mode

Mode =2.8 , 3.8, 4.7, 4.8 and 5.0

d) Standard deviation :

S=\sqrt{\sum_{i=1}^{i=20}}(x-\mu )^{^{2}} / \sqrt{n-1}

x (x-μ) (x-μ)^2
1 4.2 0.2 0.04
2 4.7 0.7 0.49
3 4.7 0.7 0.49
4 5 1 1
5 3.8 -0.2 0.04
6 3.6 -0.4 0.16
7 3 -1 1
8 5.1 1.1 1.21
9 3.1 -0.9 0.81
10 3.8 -0.2 0.04
11 4.8 0.8 0.64
12 4 0 0
13 5.2 1.2 1.44
14 4.3 0.3 0.09
15 2.8 -1.2 1.44
16 2 -2 4
17 2.8 -1.2 1.44
18 3.3 -0.7 0.49
19 4.8 0.8 0.64
20 5 1 1
16.46

standard deviation =sqrt(16.46 ) / sqrt (20-1)

                          =4.057093 / 4.35889

standard deviation =0.931